lemon1boy1:

An observer (O) is located 500 feet from a school (S). The observer notices a bird (B) flying at a 39° angle of elevation from his line of sight. How high is the bird flying over the school? You must show all work and calculations to receive full credit. A right triangle is shown with one leg measuring 500 and another leg measuring h, with an angle across from it measuring 39 degrees.

snowflake0531:



snowflake0531:

\[sin(\theta)=\frac{opposite}{hypotenuse}\\cos(\theta)=\frac{adjacent}{hypotenuse}\\tan(\theta)=\frac{opposite}{adjacent}\]

snowflake0531:

looking from the angle, you have adjacent and opposite, we would use tangent

snowflake0531:

tan(39) = x/500 x = tan(39) * 500